840:153g:Projects/project1/2008/10/09

From OpenWetWare
Jump to: navigation, search
Owwnotebook icon.png A Flower of a Different Color by "Emblazon" Report.pngMain project page
Resultset previous.pngPrevious entry      Next entryResultset next.png

Plasmid, RNA and DNA Electrophoresis

A Flower of A Different Color

Angie, Diwash, Sushma and Binu, were working on problem solving for the plasmid extraction. Angie and Diwash worked on the plasmids that were extracted on September 7th, which was the restriction digestion with EcoR1 and electrophoresis for the products to be tested. Sushma and Binu did the plasmid extraction with the cultures that were prepared September 8th (from MIT). The results from the electrophoresis gel for the plasmid showed that the bands formed on the gel were not of the expected band size. The conclusions from this experiment showed that there are problems with sterilization possibly due to the presence of RNAse. Other results from this experiment were from the pBluescript also showing incorrect band size. This could be due to step-up issues while performing the experiment or the actual protocol used for the experiment preformed. Axel and the group members working with the plasmid decided that the results from the two experiments could have had poor results due to the EcoRI used. For next week we will digest the plasmids (the third set and the one from today’s extraction) with 3 restriction enzymes. The restriction enzymes that we will be using are the new EcoR1, Spe1 and Xba1. Spe1 and Xba1 will also be used to insert the gene into plasmid.

Melissa and Rajiv preformed electrophoresis on the DNA PCR product of the two DNA leaf extractions. The band lengths that formed were around 2000 base pairs. The sequence is around 1529 base pairs according the NCBI website which concludes that there is possibility of the entire sequence being amplified with the primers used. Also due to introns within the sequence that have amplified. Other results from the gel showed that the ideal temperature for the two DNA stocks annealing temperature is 53.7C. And the ideal for PCR set-up for the two DNA stock samples is using 8.4micro-litters of DNA stock solution. Next week we will discuss with Axel the results and send the rest of the DNA PCR product to be sequenced. The sequencing will show us if the product amplified was the ideal gene sequence that we would like to work with and prepare to insert the gene sequence into the plasmid vector. If the sequence did not amplify then the primer sequences used were not specific enough and need to be redesigned.


Tiffany ran the Formeldehyde Agarose Gel Electrophoresis on the extracted RNA again today. The results again were that there was no RNA in the sample. This could possibly be due to not using RNA-ase free materials and equipment during the extraction process. We will discuss these results with Axel on Tuesday to determine our next step.